OpenStudy (anonymous):

Guys I really need help with this one!! Dx Consider the functions f(x) = 6x + 2 and g(x) = 9x - 23. What is the value of ( f over g )(4)? Please show all of your work for full credit.

OpenStudy (mertsj):

6(4)+2/9(4)-23= 26/13=2

OpenStudy (anonymous):

okay thank you! :DD

OpenStudy (anonymous):

And this one too? Consider the functions f(x) = 3x2 and g(x) = 2(x + 9). Find f[g(-7)]. Please show all of your work for full credit.